# Interpreting CLIP with Sparse Linear

### v1

@@ 4 Method
In this section we introduce SpLiCE , a method for expressing CLIP’s image representations as sparse, nonnegative, linear combinations of a concept dictionary. In the subsections that follow, we outline our design choices, including how we choose the concept dictionary in Section 4.1, how we bridge the modality gap between CLIP’s images and text representations in Section 4.2, and our method for decomposition in Section 4.3.
4.1 Concept Vocabulary
Natural language is an intuitive, interpretable, and compact medium for communicating semantic information. Thus, we choose to represent the semantic content contained in CLIP embeddings as combinations of semantic concepts expressed in natural language. We formalize this as dictionary learning problem over a concept vocabulary of natural language embeddings, where we define concepts as a semantic visual unit that can be expressed by a single English word.
Previous works have found that existing concept-based explanation methods are highly sensitive to the concept vocabulary chosen Ramaswamy et al. (2022), and a prevalent critique of methods like traditional Concept Bottleneck Models is that they rely on a well-specified and expert-chosen concept set for the task at hand. Furthermore, there often exists a direct trade-off between the size of the concept set and the interpretability offered by the method, as an explanation with too many concepts fails to be comprehensible to humans Ramaswamy et al. (2022). However, because we enforce sparsity in our concept decompositions, we can choose to have a large, overcomplete vocabulary that well spans the concept space of CLIP and thus is expert- and task-agnostic. To construct our vocabulary, we consider all one-word English tokens in the text captions of the LAION-400m dataset Schuhmann et al. (2021). We filter the captions to remove any NSFW samples and choose the top 10000 most common words to be our concept vocabulary.
4.2 Modality Alignment
In order to decompose images into text concepts, we must ensure that the representations of images span those of our concept set. However, Liang et al. (2022) show the existence of a modality gap in CLIP, where image and text embeddings can lie in non-overlapping spaces on the unit sphere. Empirically, we find that CLIP image and text embeddings exist on two cones, as the distribution of pairwise cosine similarities between MSCOCO image and text embeddings concentrate at a positive value (Fig. 3). To rectify this, we mean-center CLIP images with the image cone mean, estimated over MSCOCO ($\mathbf{\mu_{\textrm{img}}}$), and compute decompositions over the mean-centered concept vocabulary ($\mathbf{\mu_{\textrm{con}}}$). Note that the embeddings need to be re-normalized after centering to ensure they lie on the unit-sphere. If we wish to convert our decompositions back into dense representations ($\hat{\mathbf{z}}^{\textrm{img}}$), we uncenter the normalized dense embeddings $\hat{\mathbf{z}}^{\textrm{img}}$ by adding the image mean back in and normalizing once again, to ensure they lie on the same cone as the original CLIP embeddings ($\mathbf{z}^{\textrm{img}}$).
@@ 4.3 Sparse Nonnegative Concept Decomposition
We consider two desiderata when semantically decomposing dense CLIP representations: sparsity, and nonnegativity. A large body of work has shown that sparsity and nonnegativity yield greater interpretability in representations, as sparse sets can be comprehended all at once, while negative semantics remain unintuitive and difficult to conceptualize. Ramaswamy et al. (2022); Faruqui et al. (2015); ACichocki et al. (2009); Murphy et al. (2012); Fyshe et al. (2014, 2015); Lipton (2017); Ribeiro et al. (2016).
[…]
$\mathbf{C}=\left[\sigma(g(\mathbf{x}^{\textrm{con}}_{1})-\mathbf{\mu_{\textrm{con}}}),\cdots,\sigma(g(\mathbf{x}^{\textrm{con}}_{c})-\mathbf{\mu_{\textrm{con}}})\right],$
and a centered CLIP image embedding
$\mathbf{z}=\sigma(\mathbf{z}^{\textrm{img}}-\mathbf{\mu_{\textrm{img}}}),$
we seek to find the sparsest solution that gives us a cosine similarity score of at least $1-\epsilon$ for some small $\epsilon$:
@@
As is standard practice, we relax the $\ell_{0}$ constraint and reformulate this as a minimization of MSE with an $\ell_{1}$ penalty, to construct the following convex relaxation [^2] For more discussion on the relationship between Eq. (1) and Eq. (2), see Appendix, Sec. A of Eq. (1):
@@
Note that the linearity of $\mathbf{w}$ allows us to interpret the magnitude of a coordinate $\mathbf{w}_{i}$ as the prominence of concept $i$ in the image $\mathbf{x}^{\text{img}}$. With the solution to our optimization problem defined as $\mathbf{w}^{*}$, our reconstructed embedding is then
$\hat{\mathbf{z}}^{\textrm{img}}=\sigma(\mathbf{C}\mathbf{w}^{*}+\mathbf{\mu_{\textrm{img}}}).$ (3)
@@ 5 Experiments
Figure 4: Performance of SpLiCE decomposition representations on zero-shot classification tasks (bottom row) and cosine similarity between CLIP embeddings and SpLiCE embeddings (top row). We find that the semantic dictionary collected from LAION (yellow) performs the best for zero-shot classification tasks, but that the dictionary learned over images (green) is best for reconstructing embeddings with minimal cosine error.
In this section we evaluate our method to ensure that the added interpretability results in minimal loss in performance for downstream tasks and accurately reflects the semantic content of representations.
@@ 5.1 Setup
Models. All experiments shown in the main paper are done with the OpenCLIP ViT-B/32 model Ilharco et al. (2021) with results for an additional model in Appendix 13. For all zero-shot classification tasks we use the prompt template “A photo of a {}”. Datasets. We use CIFAR100 Krizhevsky et al. (2009), MIT States Isola et al. (2015), CelebA Liu et al. (2015), MSCOCO Lin et al. (2014), and ImageNetVal Deng et al. (2009) for our experiments.
Decomposition. For all experiments involving concept decomposition, we use sklearn’s Pedregosa et al. (2011) Lasso solver with a non-negativity flag and an $l_{1}$ penalty that results in solutions with $l_{0}$ norms of 15-30 (around 0.2 for most datasets), unless stated otherwise, in accordance with the threshold of 32 concepts found by Ramaswamy et al. (2022). We use a concept vocabulary chosen from a subset of LAION tokens as described in Section 4.1. Both images and concepts are centered as mentioned in Section 4.2, with the image mean used for centering computed over the MSCOCO train set and the concept mean computed over our chosen vocabulary.
@@ 5.2 Sparsity-Performance Tradeoffs
We assess the performance of SpLiCE decompositions by evaluating the reconstruction error between SpLiCE representations and CLIP embeddings, the zero-shot performance of SpLiCE , and the retrieval performance of the various embeddings. We compare the performance of decompositions generated from our semantic concept vocabulary to decompositions over a random vocabulary and dictionary learned vocabularies. All vocabularies include 10,000 concepts, where the random vocabulary is sampled from a Gaussian distribution and normalized to lie on a hypersphere. The learned vocabularies are generated by using the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) Beck and Teboulle (2009) to learn the optimal dictionary over either the MSCOCO image corpus or caption corpus.
In Figure 4, we visualize the cosine reconstruction and zero-shot accuracy of image decompositions with various dictionaries. We find that while our semantic concept dictionary does not reconstruct samples as well as the learned dictionaries with regard to cosine similarity, it significantly outperforms the decompositions with random and learned dictionaries for zero-shot classification tasks. In particular, decompositions with low sparsity ($<$5 nonzero concepts) still provide enough semantic information to classify CIFAR100 and MIT States images with similar performance to the vanilla CLIP embeddings. We note that performance on ImageNet is worse than the other datasets because many ImageNet classes are animal species that cannot easily be described by a few single-word semantic concepts (e.g. ‘boston bull’). Adding these class labels to our concept dictionary increases performance significantly, as shown by the dotted yellow line in Figure 4. We note that interestingly, concept dictionaries learned over images outperform dictionaries learned over text for cosine reconstruction, but the opposite is true for zero-shot performance at high sparsities. We evaluate probing performance in the Appendix (Tables 3, 4).
We finally test the performance of SpLiCE embeddings on text-to-image and image-to-text retrieval tasks. We evaluate retrieval over various 1024 sample subsets of MSCOCO, and assess recall performance for the top-k closest embeddings of the opposite modality. We visualize $k=\{5\}$ for text-to-image retrieval in Figure 5, with $k=\{1,10\}$ and image-to-text in the Appendix (Figure 12). We note that while the decompositions do not perfectly match the performance of dense CLIP embeddings at high sparsities, an order of magnitude reduction in memory ($l_{0}$ norm $\approx 50$) only leads to a drop in recall of about 15%.
5.3 Concept Based Explanations of Images
We qualitatively evaluate SpLiCE decompositions as measures of the semantic content of the images they represent. In Figure 2 we provide six sample decompositions from MSCOCO with their corresponding captions. We display the top six concepts for each image, and find that they generally well-describe the semantics of the images. As such, we propose that these decompositions can be used as image taggers, or simply as concept-based explanations of the semantic content of CLIP embeddings when used in downstream tasks such as classification.
Figure 5: Performance of SpLiCE representations for text-to-image retrieval on MSCOCO.
5.4 Concept Based Explanations of Classes or Datasets
While concept decompositions can be highly useful for explaining the predictions and decisions of any systems relying on CLIP embeddings, we also propose that this method can be used to better understand and summarize datasets. By decomposing entire classes or whole datasets, we can find the underlying semantic distribution of the data. We visualize the top 7 most common concepts in three ImageNet classes (‘Bakery’, ‘Balloon’, ‘Indian elephant’, and ‘African elephant’) in Figure 6. These decompositions yielded interesting insights into the classes themselves; the second most common concept in the ‘Bakery’ class was “cupcake,” and we indeed found that 7/50 photos in the test class were close-up photos of cupcakes. The ‘Balloon’ class does not feature blow up party balloons, but rather hot-air balloons, as indicated by the concepts “parachute” and “flying”. For the two elephant classes, we see that the decompositions contain information about the backgrounds, such as “jungle”, “kerala”, and “africa”, thus encoding the nuances between the two species.
Figure 6: SpLiCE decompositions of ImageNet ‘Bakery’, ‘Balloon’, ‘Indian Elephant’, ‘African Elephant’ classes.
@@ 6 Case Studies and Applications of SpLiCE
In this section, we present three case studies using SpLiCE : (1) spurious correlation and bias detection in datasets 6.1, (2) debiasing classification models 6.2, and (3) monitoring distribution shift 6.3.
6.1 Case Study: Spurious Correlation Detection
A key motivating application for SpLiCE is the ability to automatically summarize datasets semantically. We find that this is particularly helpful for auditing bias or spurious correlations in datasets, which would otherwise require labels of both the subgroup and the spuriously correlated feature. We present a case study of the two adult classes of CIFAR100: ‘man’ and ‘woman’. Upon decomposing these classes, we found that “bra” and “swimwear” were two of the top ten most common concepts in the ‘woman’ class. On the other hand, the only clothing-related concepts that appear in the top 50 most activated concepts for ‘men’ are {“uniform”, “tuxedo”, “apparel”}. We visualize a histogram of the concept weights on swimwear and undergarment related concepts {“swimwear”, “bra”, “trunks”, “underwear”} in Figure 7 across both the train and test sets, and find that these concepts are much more likely to be activated for women than men. This phenomenon was confirmed via manual inspection of the CIFAR100 dataset, where at least 8 of the 100 images in the ‘woman’ test class feature women in bikinis, revealing stereotype bias in this popular dataset.
Figure 7: Distribution of “Swimwear” concept in ‘Woman’ and ‘Man’ classes of CIFAR100.
6.2 Case Study: Model Editing
Another key application of concept decompositions is using them to isolate specific information, allowing for fine-grained intervention. This intervention can be performed on the embeddings themselves, where we can zero out a concept and reconstruct the input without that information, or probes built upon the decompositions, where we can manually alter the relationships between concepts and class labels (similar to a concept bottleneck Koh et al. (2020)).
Consider a facial recognition classifier trained on images of CelebA. We wish for this classifier to be invariant to the presence of glasses, but we are unsure if there exist spurious correlations between certain individuals and the presence of eyewear. We thus wish to remove the concept of “eyewear” or “glasses” from the image representations, such that they cannot be leveraged for downstream classification. We can do so by simply zero-ing out any weight placed upon those concepts in our SpLiCE decompositions, and evaluate if a zero-shot classifier can detect the presence of glasses. We report performance of zero-shot classifier in identifying the concept “glasses” before and after intervention in rows ‘ZS SpLiCE ’ and ‘Intervention SpLiCE ’ respectively. Vanilla CLIP performance is given as a baseline in ‘ZS CLIP’. We see that removing that intervention reduces accuracy of the zero-shot classifier from 0.88 to 0.69 (where 0.50 is random). We also evaluate intervention on linear probes, where we ablate probe weights of the concept, and see that the glasses probe’s accuracy drops from 0.88 to 0.59 after invention. We provide gender classification performance for the same as a proxy for facial recognition, and see that intervention on “glasses” for both zero-shot and probing tasks has minimal effect on gender classification.
Table 1: Evaluation of Intervention on the Concept ‘Glasses’
@@ 6.2 Case Study: Model Editing
6.3 Case Study: Distribution Shift Monitoring
We present a final case study using SpLiCE to monitor distribution shift. This can help identify differences between training and inference distributions or evaluate how a continually sampled dataset changes over time. In this experiment we consider the Stanford Cars dataset, which contains photos of cars from 1991 to 2012, including their make and year labels. By decomposing photos of cars from each year, we can view how the distribution changed yearly. We visualize the weights of the concepts “convertible” and “yellow” from our decompositions, as well as the actual percentage of cars from each year that were convertibles or yellow in Figure 8. Note the right-hand y-axis, corresponding to the weight of the given concept $c_{i}$ over the sum of the weights of all concepts $\sum_{i}c_{i}$, does not have a meaningful unit of measure or scale. We find that the trends in the groundtruth concept prevalence generally closely match that of the predicted/decomposed concepts, allowing us to visualize which years convertibles or yellow cars were popular or out-of-distribution with respect to other years.

### v2

@@ 4 Method
In this section, we introduce SpLiCE , a method for expressing CLIP’s image representations as sparse, nonnegative, linear combinations of concept dictionary elements. We begin by framing this problem as one of sparse recovery. We then discuss our design choices, including how we choose the concept dictionary and how to address the modality gap between CLIP’s images and text representations. Finally, we formalize the optimization problem used in this work.
@@ 4.1 Sparse Nonnegative Concept Decomposition
Our goal is to construct decompositions of dense CLIP representations that are human-interpretable, useful, and faithful. To do so, we formulate decomposition as a sparse recovery problem with three main desiderata. First, for the decompositions to be interpretable to humans they must be comprised of human interpretable atoms. We argue that language is a naturally interpretable interface for humans, and construct our concept vocabulary $\mathbf{C}$ out of 1- and 2-word atoms, such as “coffee”, “silver”, and “birthday party”. Second, our decompositions must be simple and concise, which can be formulated as a sparsity constraint on the recovery. A large body of work in computational linguistics [42, 43, 44, 14], neuroscience [45, 46], and interpretability [47, 48, 30] have demonstrated that a human-aligned semantic model should be sparse in representation. Furthermore, [47] found that users can best understand explanations with fewer than 32 concepts while in linguistics, [49, 50, 51] find participants describe concepts and objects with up to 20 semantic properties, motivating our desiderata of sparsity. Third, our decompositions must be constructive, i.e., we must decompose representations in terms of their constituent concepts. For this reason, we require the weights of decompositions to be strictly nonnegative, to avoid having “negative” concept weights which do not always carry semantic meaning. Furthermore, prior work by Zhou et al. [30] has argued that “negations of concepts are not as interpretable as positive concepts.” More specifically, while a small set of concepts have well-defined antonyms which may be viewed as their negative counterparts (‘‘day’’ $\leftrightarrow$ ‘‘night’’), negative concepts do not carry semantic meaning in general (‘‘tiger’’ $\leftrightarrow$ ??). Furthermore, we find that even when antonyms exist, they are not negatives of each other in CLIP latent space (see Appendix B.10). To avoid dependence on negative weights and ensure that all concepts are captured, we construct an overcomplete dictionary containing a wide range of concepts, including antonyms. We build on top of this literature and provide a semantic decomposition satisfying these properties suitable for multimodal models like CLIP.
Figure 3: Performance of SpLiCE decomposition representations on zero-shot classification tasks (bottom row) and cosine similarity between CLIP embeddings and SpLiCE embeddings (top row). Our proposed semantic dictionary (yellow) closely approximates CLIP on zero-shot classification accuracy, but not on the cosine similarity. This indicates that SpLiCE captures the semantic information in CLIP, but not its non-semantic components, explaining both the high zero-shot accuracy and low cosine similarity. See §5.2 for discussion.
Concept Vocabulary. Natural language is an intuitive, interpretable, and compact medium for communicating semantic information. Thus, we choose to represent the semantic content contained in CLIP embeddings as combinations of natural language semantic concepts, where we define concepts as semantic units that can be expressed concisely, by one- or two-word phrases. Given that CLIP is used in a wide variety of downstream applications and is trained without a specific task in mind, we want our concept dictionary to be task-agnostic and to span all possible concepts CLIP could have learnt. To construct this vocabulary, we consider the most frequent one- and two-word bigrams in the text captions of the LAION-400m dataset [52], the dataset that most CLIP variants are trained on. We filter the captions to remove any NSFW samples and prune our concept set such that no two concept embeddings have a cosine similarity greater than 0.9. We also remove bigrams highly similar ($>0.9$ cosine similarity) to the average of their individual words. We finally choose the top $10,000$ most common single-word concepts and the top $5000$ most common two-word concepts as our concept vocabulary. We note that this vocabulary offers distinct advantages over those used in prior works. In particular, it is task-agnostic, meaning that the efficacy of the decomposition is (in principle) independent of individual datasets. Furthermore, this dataset imposes minimal priors from outside curators, such as human experts or LLMs [22, 23, 24]. This allows us to interpret data through the lens of CLIP, to understand the information encoded, including potential biases and mistakes.
Modality Alignment. In order to decompose images into nonnegative combinations of text concepts, we must ensure that our concept set spans the space of possible image embeddings. However, [53] show the existence of a modality gap in CLIP, where image and text embeddings can lie in non-identical spaces on the unit sphere. We empirically find that CLIP image and text embeddings exist on two cones, as the distribution of pairwise cosine similarities between pairs of MSCOCO images and pairs of MSCOCO text captions concentrate at positive values, whereas the distribution of pairwise cosine similarities across modalities concentrates closer to zero. (See Appendix Fig. 7). Not only does this prevent nonnegative decomposition, it also violates Assumption 4 from Section 3. To rectify this, we mean-center CLIP images with the image cone mean, estimated over MSCOCO ($\mathbf{\mu_{img}}$), and compute decompositions over the mean-centered concept vocabulary ($\mathbf{\mu_{con}}$). Note that the embeddings need to be re-normalized after centering to ensure they lie on the unit sphere. To convert our decompositions back into dense representations ($\hat{\mathbf{z}}^{\text{img}}$), we uncenter the normalized dense embeddings $\hat{\mathbf{z}}^{\text{img}}$ by adding the image mean back in and normalizing once again, to ensure they lie on the same cone as the original CLIP embeddings ($\mathbf{z}^{\text{img}}$).
Optimization Problem. Our optimization problem is formulated as follows. Let $\sigma(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$ be the normalization operation. Given a set of semantic concepts $\mathbf{x}^{\textrm{con}}=~{}$[“dog”, “tabby cat”, “cloudy”, $\cdots$ ], we construct a centered vocabulary $\mathbf{C}=\left[\sigma(g(\mathbf{x}^{\textrm{con}}_{1})-\mathbf{\mu_{\textrm{con}}}),\cdots,\sigma(g(\mathbf{x}^{\textrm{con}}_{c})-\mathbf{\mu_{\textrm{con}}})\right]$, where we recall that $g(\cdot)$ is the CLIP text encoder. Now, given the dictionary $\mathbf{C}$ and a centered CLIP embedding $\mathbf{z}=\sigma(\mathbf{z}^{\textrm{img}}-\mathbf{\mu_{\textrm{img}}})$, we seek to find the sparsest solution that gives us a cosine similarity score of at least $1-\epsilon$ for some small $\epsilon$:
@@
As is standard practice, we relax the $\ell_{0}$ constraint and reformulate this as a minimization of MSE with an $\ell_{1}$ penalty, to construct the following convex relaxation [^3] For more discussion on the relationship between Eq. (1) and Eq. (2), see Appendix, Sec. A.2 of Eq. (1):
@@
Given the solution to the above problem $\mathbf{w}^{*}$, our reconstructed embedding is: $\hat{\mathbf{z}}^{\textrm{img}}=\sigma(\mathbf{C}\mathbf{w}^{*}+\mathbf{\mu_{\textrm{img}}}).$
@@ 5 Experiments
In this section, we evaluate our method to ensure that SpLiCE decompositions are interpretable, performant, and accurately reflect the semantic content of representations.
@@ 5.1 Setup
Models. All experiments shown in the main paper are done with the OpenCLIP ViT-B/32 model [54] with results for an additional model in Appendix B.14. For all zero-shot classification tasks, we use the prompt template “A photo of a {}”. Datasets. We use CIFAR100 [55], MIT States [56], CelebA [57], MSCOCO [58], and ImageNetVal [59] for our experiments with results for additional datasets in the Appendix (Section B.4)
Decomposition. For all experiments involving concept decomposition, we use sklearn’s [60] Lasso solver with a non-negativity flag and an $l_{1}$ penalty that results in solutions with $l_{0}$ norms of 5-20 (around 0.2-0.3 for most datasets).We use a concept vocabulary chosen from a subset of LAION tokens as described in Section 3. Both image embeddings and dictionary concepts are centered and normalized as mentioned in Section 4.1, with the image mean used for centering computed over the MSCOCO train set and the concept mean computed over our chosen vocabulary.
@@ 5.2 Sparsity-Performance Tradeoffs
We assess the performance of SpLiCE decompositions by evaluating the reconstruction error in terms of cosine similarity between SpLiCE representations and CLIP embeddings, the zero-shot performance of SpLiCE  decompositions, and the retrieval performance of SpLiCE embeddings. We compare the performance of decompositions generated from our semantic concept vocabulary to decompositions over random vocabulary and learned dictionary vocabulary baselines. All vocabularies are of size 15,000 concepts. The random vocabulary is sampled from a 512-dimensional normalized Gaussian distribution. The learned vocabularies are generated by using the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) [61] to learn optimal dictionaries given our sparse recovery problem (optimizing Equation (1) for both $\mathbf{C}$ and $w$). Note that we learn separate dictionaries $\mathbf{C}_{\text{img}}$ and $\mathbf{C}_{\text{text}}$ to reconstruct MSCOCO image and text embeddings respectively. In Figure 3, we plot the cosine reconstruction and zero-shot accuracy of image decompositions with the various dictionaries. We evaluate probing performance (Tables 3, 4) and text-to-image and image-to-text retrieval in the Appendix (Figure B.3).
These results overall show SpLiCE efficiently navigates the interpretability-accuracy Pareto frontier and retains much of the performance of black-box CLIP representations with the semantic, human-interpretable LAION dictionary, significantly outperforming other dictionaries on semantic tasks such as zero-shot classification, probing, and retrieval. At the same time, we find that our semantic LAION dictionary does not result in accurate cosine similarity reconstructions of the original CLIP, often being on par with using random dictionaries. We believe this is because CLIP encodes both semantics of the underlying image and non-semantic "noise", which violates Assumption #2 in Section 3. Given that our SpLiCE decompositions only aim to encode semantics, they are unable to encode non-semantic aspects in the underlying representation, thus causing poor alignment in the cosine similarity sense, while simultaneously exhibiting excellent alignment on semantic tasks such as zero-shot accuracy. For ImageNet, we find that many classes are animal species that cannot easily be described by 1-2 words (e.g. ‘red-breasted merganser’, ‘American Staffordshire terrier’). Adding these class labels to our concept dictionary increases performance significantly, as shown by the dotted yellow line in Figure 3.
Figure 4: Left: SpLiCE decompositions of ImageNet ‘African Elephant’, ‘Curly-coated Retriever’, ‘Monarch Butterfly’, ‘Digital Clock’ classes. Right: Distribution of “Swimwear" concept in ‘Woman’ and ‘Man’ classes of CIFAR100.
5.3 Ablation Studies
We perform ablation studies to evaluate the effectiveness of the design decisions of SpLiCE , including the choice of vocabulary, the nonnegativity of the decompositions, and the modality alignment by ablating each choice and observing the effect on three metrics: zero-shot accuracy on CIFAR100, cosine similarity between the reconstructed and original embeddings of CIFAR100, and semantic relevance on MSCOCO. The first two metrics are the same as those presented in Figure 3. We compute semantic relevance by tokenizing and filtering stop-words from the MSCOCO human-generated captions and embedding each token with CLIP. Then, we take all non-zero concepts output by SpLiCE and compute the Hausdorff distance between the sets of SpLiCE concepts and caption token embeddings. This essentially measures how aligned decompositions are with human captions. We observe that replacing our dictionary with the LLM-generated concept dictionary used by [22, 24, 23] significantly worsens the decomposition in terms of zero shot accuracy and cosine reconstruction. While allowing for negative concept weights improves cosine reconstruction marginally, it decreases the semantic relevance of the decompositions, as negative concepts frequently correspond to concepts not present in images, and as such, are unlikely to be represented by human captions. Finally, we see that modality alignment is necessary across all three metrics. Overall, these ablation studies show that each aspect of SpLiCE is necessary for creating human-interpretable, semantically relevant and highly performant decompositions.
Figure 5: Ablation study evaluating the efficacy of SpLiCE design choices across three metrics: Zero-shot accuracy, cosine reconstruction, and semantic relevance of recovered tags. We find that all of our design choices, namely non-negativity, modality alignment, and usage of large task-agnostic dictionary are essential to performance. See §5.3 for discussion.
5.4 Qualitative Assessment of Decompositions
Concept Decompositions for Images. We visualize SpLiCE decompositions to qualitatively assess the semantic content of the images they represent. In Figure 2 we provide six sample decompositions from MSCOCO with their corresponding captions. We display the top seven concepts for each image and find that they generally well describe the semantics of the images. We also find that these qualitative examples yield interesting and unexpected insights into both CLIP and the data. In the top left image, we see that the decomposition includes the text present on the sign in the image, revealing that CLIP prioritizes text in images over objects. For the bottom left image, the decomposition correctly includes the concept “macro shot”, revealing that CLIP encodes information regarding geometric perspective. The bottom right decomposition similarly features the concept “blackandwhite bw”, indicating that CLIP encodes not only the objects present in images but also information about the lighting and color. Overall, these results suggest that SpLiCE may also be used as a zero-shot image tagging method to understand images.
Concept Histograms for Datasets. Beyond concept-based explanations of individual images, we propose that SpLiCE can be used to better understand and summarize collections of images, such as entire datasets. To compute concept decompositions of sets of images, we decompose each individual image and aggregate the results, which we use to generate concept histograms of the dataset. We visualize four concept histograms for the ImageNet classes ‘African Elephant’, ‘Curly-coated Retriever’, ‘Monarch Butterfly’, and ‘Digital Clock’, in Figure 4. These decompositions provide information about the distribution of the data and how CLIP represents it. For example, digital clocks are differentiated from analog clocks through the concepts “lcd display” and “countdown”. Monarch butterflies are highly correlated with the concept “lilac flowers" in ImageNet, which we validated through manual inspection (nearly half of the monarch butterfly images in the validation set feature purple flowers). Interestingly, ‘Curly-coated retrievers’ are represented as combinations of “poodle”, “retriever dog”, and “black dog”, which perfectly describe the main characteristics of them: black retrievers with poodle-textured fur.
@@ 6 Case Studies and Applications of SpLiCE
In this section, we present two example case studies using SpLiCE : (1) spurious correlation and bias detection in datasets and (2) debiasing classification models. We present additional case studies for (1) and (2), as well as (3) monitoring distribution shift in Appendix B.6, B.7, B.8 B.9. We also present results from a user study to evaluate the human interpretability of SpLiCE in Appendix B.1, where we find that users prefer explanations generated by SpLiCE over existing Concept Bottleneck Model-based methods.
Discovering Spurious Correlations in CIFAR100.
Existing methods to detect spurious correlations in datasets generally require subgroup and attribute labels or rely on manual human inspection of images (see [62] for an overview), making it hard to scale to large datasets. SpLiCE , on the other hand, allows for fast automatic detection of such biases, without any labels, training, or even a task. To illustrate this, we study two classes of CIFAR100: ‘man’ and ‘woman’, in Figure 4. Upon decomposing these classes, we found that {“bra”,“swimwear” } were two of the top ten most common concepts in the ‘woman’ class. On the other hand, the only clothing-related concepts that appear in the top 50 most activated concepts for ‘man’ are {“uniform”, “tuxedo”, “apparel”}. We visualize a histogram of the concept weights on swimwear- and undergarment-related concepts {“swimwear”, “bra”, “trunks”, “underwear”} across both the train and test sets, and find that these concepts are much more likely to be activated for women than men. Manual inspection of CIFAR100 verifies the trend highlighted by SpLiCE, where at least 70 of the 600 images in the ‘woman’ class feature women in bikinis, underclothes, or even partially undressed, revealing stereotype bias in this popular dataset. We provide a similar study of the concept “desert” with respect to the ‘camel’ and ‘kangaroo’ classes in CIFAR100 in Appendix B.6.
Model Editing on CelebA Attribute Classifiers.
Table 2: Evaluation of intervention on the concept ‘Glasses’ for the CelebA dataset. SpLiCE allows for surgical removal of information related to whether or not someone is wearing glasses, without impacting other features such as gender. (ZS = Zero Shot Accuracy)
@@
Concept-based representations unlock a key application: being able to intervene on and edit models. This edit can be performed in two equivalent ways: either on the concept representations themselves, where we can zero out a concept or on linear probes built upon the decompositions, where we can edit the weight matrix between concepts and class labels (similar to concept bottleneck models [18]). Here, we evaluate the efficacy of SpLiCE for these forms of model editing. Specifically, we consider two tasks on CelebA, classifying gender and whether the subject is wearing glasses. To test representation editing, we remove the concept of “eyewear” or “glasses” from CelebA image representations by zeroing out any weight placed on these concepts in our SpLiCE decompositions and evaluate classifier performance. We report the performance of zero-shot classification and linear probes over our SpLiCE representation in Table 2. In both cases, we find that we can surgically target and remove information pertaining to glasses and reduce classifier performance while preserving information relevant to gender classification. We perform a similar experiment on the Waterbirds dataset [63] to remove spurious background signals in B.7.